The magnetization within a bar of some metal alloy is 1.2×106 A/m when the H field is 200 A/m.

The magnetic susceptibility χm of this alloy is,
a, 4000
b, 4560
c, 5400
d, 6000
1
Expert's answer
2021-10-06T15:53:39-0400

Magnetic susceptibility of the alloy can be found as follows:


"\chi_m=\dfrac{M}{H}=\dfrac{1.2\cdot10^6\ \dfrac{A}{m}}{200\ \dfrac{A}{m}}=6000."

Answer:

d. 6000